# R-CoT[Uncaptioned image]: Reverse Chain-of-Thought Problem Generation for Geometric Reasoning in Large Multimodal Models

### v1

@@ 2 Related Work
Recent research aimed at improving geometric reasoning in LMMs can be broadly divided into two categories. The first category focuses on inspiring geometric reasoning ability during the inference stage, while the other attempts to improve reasoning ability through targeted training.
Inspiring Model Potential During Geometric Inference.
@@
For the inference process,  Zhao et al. (2024a) employs the chain of thought in visual and symbolic language modes to cross-validate and correct each other for the final result.  Hu et al. (2024) utilizes code to generate images and solve problems through a visual chain of thought. Meanwhile,  Mouselinos et al. (2024) uses a LLM as an agent to call external tools.
Improving Model Reasoning Ability During Geometric Training.
@@
For model training, symbolic geometry solvers like GeoS  (Seo et al., 2015), Inter-GPS  (Lu et al., 2021), and S2G  (Tsai et al., 2021) aim to build formal language systems that use formal language for deductive reasoning on geometry problems. These systems are manually designed for formal languages with relatively small datasets, e.g. the GeoS dataset containing 186 problems and the Geometry3k  (Lu et al., 2021) dataset containing about 3000 problems. The size of the datasets has increased slightly with the advent of neural geometric solvers, such as UniGeo  (Chen et al., 2022), GeoQA  (Chen et al., 2021), GeoQA+  (Cao & Xiao, 2022), and PGPS9K  (Zhang et al., 2023a), with a total size of around 25k. The above datasets are collected manually, with high labeling costs and limited scale. With the rise of LMMs, these data scales are far from satisfactory for training, so many methods are devoted to building larger datasets. G-LLaVA  (Gao et al., 2023) uses an LLM to reword original Q&A pairs in the GeoQA and Geometry3k dataset, resulting in 115k geometric Q&A data and 60k alignment data but does not increase the diversity of images and knowledge points. GeomVerse  (Kazemi et al., 2023) uses a code-written engine to generate accurate geometric images and Q&A pairs, but there is still a certain gap between the generated images and real-world geometric images. Additionally, the questions generated by the template lack diversity.
To synthesize geometry data with accuracy and diversity, we introduce R-CoT, a novel geometry data generation pipeline that addresses visual hallucinations and reasoning limitations in LMMs. This pipeline effectively generates the GeoMM dataset, featuring high-fidelity geometric images with accurate and diverse Q&A pairs.
@@ 4.3 Ablation Study
Figure 7: Compared with existing datasets at different data scales.

### v2

@@ 2 Related Work
Inspiring Potential in Geometry Inference.
@@
For the inference process,  Zhao et al. (2024a) employs the chain of thought in visual and symbolic language modes to cross-validate and correct each other for the final result.  Hu et al. (2024) utilizes code to generate images and solve problems through a visual chain of thought. Meanwhile,  Mouselinos et al. (2024) uses a LLM as an agent to call external tools. These methods use external assistance to enhance reasoning results without really enhancing the reasoning ability of LMMs.
Enhancing Reasoning in Geometry Training.
@@
For model training, symbolic geometry solvers like GeoS  (Seo et al., 2015), Inter-GPS  (Lu et al., 2021), and S2G  (Tsai et al., 2021) aim to build formal language systems that use formal language for deductive reasoning on geometry problems. These systems are manually designed for formal languages with relatively small datasets, e.g. the GeoS dataset containing 186 problems and the Geometry3k  (Lu et al., 2021) dataset containing about 3000 problems. The size of the datasets has increased slightly with the advent of neural geometric solvers, such as UniGeo  (Chen et al., 2022), GeoQA  (Chen et al., 2021), GeoQA+  (Cao & Xiao, 2022), and PGPS9K  (Zhang et al., 2023a), with a total size of around 25k. The above datasets are collected manually, with high labeling costs and limited scale. With the rise of LMMs, these data scales are far from satisfactory for training, so many methods are devoted to building larger datasets. G-LLaVA  (Gao et al., 2023) uses an LLM to reword original Q&A pairs in the GeoQA and Geometry3k dataset, resulting in 115k geometric Q&A data and 60k alignment data but does not increase the diversity of images and knowledge points. GeomVerse  (Kazemi et al., 2023) uses a code-written engine to generate accurate geometric images and Q&A pairs, but there is still a certain gap between the generated images and real-world geometric images. Additionally, the questions generated by the template lack diversity. To synthesize geometry data with accuracy and diversity, we introduce R-CoT, a novel geometry data generation pipeline that addresses visual hallucinations and reasoning limitations in LMMs. This pipeline effectively generates the GeoMM dataset, featuring high-fidelity geometric images with accurate and diverse Q&A pairs.
Reverse Engineering.
Recently, Reverse Engineering has gained significant interest within the large model research community  (Xue et al., 2023; Yuan et al., 2024; Jiang et al., 2024; Weng et al., 2022). Specifically,  Jiang et al. (2024); Weng et al. (2022) reverse-verifies the accuracy of the results by exchanging conditions and answers.  Xue et al. (2023) uses reverse reasoning to find missing conditions or errors, thus refining the reasoning process.  Yuan et al. (2024) uses initial reasoning to obtain model preferences and reverse-optimizes the prompt to guide the model better. The above approaches do not change the model weights to improve reasoning at the source and are inefficient for verifying and correcting complex reasoning steps. Different from their goal of optimizing the prompt followed by further reasoning or verifying the accuracy of the answer, R-CoT emphasizes generating the answer first before generating the question in reverse during the data generation process. This approach reduces the inaccuracy of answers due to overly complex questions and generates accurate geometry data for model training.
@@ 4 Experiments
Figure 7: Compared with existing datasets at different data scales.